OpenStudy (anonymous):

An object falls because of gravity at a rate of -980 cm/s2 (- sign is indicative of a downward direction) for 10. seconds. If it starts from rest (vo = 0) and its position starts from do= 0.0, where is it 10. seconds later?

OpenStudy (anonymous):

Equation 1: Displacment with Constant Acceleration. This equation is useful for finding how far an object travels under constant acceleration but may also be useful for finding the distance required for an object to reach a certain speed or to come to a stop. Equation 1: df = do + vot + ½ at2 Equation 2: Displacement with Constant Acceleration. This equation may be used to find the displacement of any object moving with constant acceleration. Equation 2: df = do + ½ (vo + vf)t Equation 3: Velocity with Constant Acceleration. This equation can be used to find the final velocity of an object moving with contant acceleration after it has accelerated at a constant rate for any time interval. Equation 3: vf = vo + at Equation 4: Final Velocity after any Displacement. Remember to take the square root of the right side of the equation to find the final velocity. The square root may be either positive or negative. If you have been consistent in your use of sign convention, you will be able to determine the value that is correct by reasoning based on the direction of motion. Equation 4: vf2 = vo2 + 2aΔd Note: If do and vo = 0, equations 1-4 simplify to the following: Equation 5: df = ½ at2 (from equation 1) Equation 6: d = ½ vft (from equation 2) Equation 7: vf = at (from equation 3) Equation 8: vf2 = 2aΔd (from equation 4)

OpenStudy (anonymous):

these are the equations and im confused

OpenStudy (anonymous):

there are "3" kinematic equations. \[1)\;v_f=v_0+at\\2)\;\Delta d=v_0t+{1\over2}at^2\\3)\;v_f^2=v_0^2+2a\,\Delta d\]

OpenStudy (anonymous):

all other forms are just derivatives of these "3"

OpenStudy (anonymous):

now, from the information that we have, which of the three suits our purpose?

OpenStudy (anonymous):

will eq. (2) serve the purpose?

OpenStudy (anonymous):

i was thinking 3....

OpenStudy (anonymous):

but i think 2 now makes sense

OpenStudy (anonymous):

to use "3", we do not have \(v_f\) and \(\Delta d\)

OpenStudy (anonymous):

all these equations need "3" information and they give you a fourth one. based on what you have and what you want, use the respective equation

OpenStudy (anonymous):

\[v_0=0{\rm cm/s}\\g=-980{\rm cm/s^2}\\t=10{\rm s}\\\Delta d=?\]